The pH of a 0.018 M aqueous solution of 3,5-dimethylpyridine (C7H9N) is 9.20. Calculate Kb. (Assume Kw = 1.00 ✕ 10−14.)
Answer #1

use:

pH = -log [H+]

9.2 = -log [H+]

[H+] = 6.31*10^-10 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(6.31*10^-10)

[OH-] = 1.585*10^-5 M

C7H9N dissociates as:

C7H9N +H20 0.018 0.018-x C7H9NH++OH

Kb = [C7H9NH+][OH-]/[C7H9N]

Kb = x*x/(c-x)

Kb = 1.585*10^-5*1.585*10^-5/(0.018-1.585*10^-5)

Kb = 1.397*10^-8

Answer: 1.40*10^-8